OpenStudy (anonymous):

Write in exponential form ^6√ b

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt[6]{b} }\) like this ?

OpenStudy (anonymous):

yes @SolomonZelman I couldn't find out how to do that symbol!

OpenStudy (anonymous):

was that supposed to be a divided by d? @SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt[A]{b^{D} }=\sqrt[A]{b^{D}} =b^{D/A}}\) \(\LARGE\color{blue}{ \bf \sqrt[6]{b}=\sqrt[6]{b^{1}} =???}\)

OpenStudy (solomonzelman):

I made a mistake before, now the correct rule is there.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt[6]{b}=\sqrt[6]{b^{1}} =b^{1/6} }\)

OpenStudy (anonymous):

Will it always be one? Or will there someetimes be a actually exponent on b? @SolomonZelman

OpenStudy (anonymous):

3√ a^2 so would this be A^3/2? @SolomonZelman

OpenStudy (solomonzelman):

, yes but the other way, 2/3 the exponent inside the root is the nominator, and the exponent next to the root, like the 3 in your case is the denominator \(\LARGE\color{blue}{ \bf \sqrt[3]{a^{2} }=\sqrt[3]{a^{2}} =a^{2/3} }\)

OpenStudy (anonymous):

okay I think I understand completely. So if it is √ y would it be y^1/2? @SolomonZelman

OpenStudy (solomonzelman):

yes, exactly !!!!

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf a^{-1/2} =?}\) can you do this one ?